# VDebugger: Harnessing Execution Feedback for Debugging Visual Programs

### v2

@@ 1 Introduction
Complex visual reasoning is a crucial yet challenging problem that often requires compositionally synthesizing multiple reasoning steps before drawing the final conclusion. For example, to answer the visual question in Figure 1: “Do the skiers wear jackets of the same color?”, one must identify all skiers, determine the colors of their jackets, and assess whether the colors are the same. End-to-end vision-language models (VLMs) excel at individual tasks such as object detection (Li et al., 2022b) and visual instruction following (Liu et al., 2023). However, they struggle to generalize to complex tasks requiring compositional reasoning and inherently lack interpretability (Surís et al., 2023; Yüksekgönül et al., 2023; Hu et al., 2023).
@@ 3 VDebugger Framework
where is a confidence score representing how likely the program is correct. Concretely, first generates a correctness token chosen from representing whether the program is correct or incorrect, so the probability assigned to token can serve as . If the token is generated, further generates the error location . Here, the location is a span within program P defined by its start and end positions, which can be a word, a line, multiple lines, or any continuous segment.
@@ 4 Training of VDebugger
and we apply mask-best sampling to at most tokens. The formal algorithm is in Alg. 2. As shown in Table 2, mask-best dramatically increases the rate at which an error is successfully injected by up to 10 times (from to ). We manually analyze and categorize 200 errors injected into GQA dataset. As shown in Figure 3, greedy sampling generates a large number of superficial errors referencing variables before their creation. In contrast, mask-best sampling produces a broader range of more complex and diverse errors.
@@ 5.1 Results
| Critique Acc | Refiner SR | Task Performance | |||||
@@
| GQA | NLVRv2 | RefCOCOg | |
| Base VLM | 44.7 | 55.4 | 50.1 |
| CodeLlama-70b as code generator | |||
| No Debugging | 46.1 | 68.1 | 53.3 |
| VDebugger (7b) | 48.8 (+2.7) | 68.9 (+0.8) | 56.7 (+3.4) |
| VDebugger (13b) | 48.9 (+2.8) | 68.6 (+0.5) | 56.4 (+3.1) |
| DeepSeek-Coder-33B as code generator | |||
| No Debugging | 47.3 | 67.9 | 55.4 |
| VDebugger (7b) | 48.9 (+1.6) | 67.8 (-0.1) | 57.1 (+1.7) |
| VDebugger (13b) | 48.9 (+1.6) | 67.4 (-0.5) | 57.7 (+2.3) |
| GPT-3.5 as code generator | |||
| No Debugging | 45.9 | 63.9 | 57.5 |
| VDebugger (7b) | 49.3 (+3.4) | 68.8 (+4.9) | 58.8 (+1.3) |
| VDebugger (13b) | 49.6 (+3.7) | 68.7 (+4.8) | 58.8 (+1.3) |
Ablation study. We investigate the contribution of each component as shown in Table 4. The experiments are performed on VDebugger fine-tuned from CodeLlama-7B. Specifically, we aim to: (1) assess the individual contributions of critic and refiner components, and (2) evaluate the benefits of execution feedback. Our results show that while the critic consistently achieves high accuracy of 67% to 80%, the refiner success rate is less reliable and dramatically varies from 10% to 57% across different datasets. When enhanced with execution feedback, the critic achieves more performance gains while the benefits to refiner performance are minimal. When reflected in the final performance on the downstream tasks, execution feedback consistently brings benefits on all datasets. In general, VDebugger can reliably perform self-critique utilizing execution feedback, and the remaining challenges mainly lie in correcting the program after the errors are identified.
Performance by iteration. VDebugger can perform iterative debugging until the critic determines the program as correct. Figure 4 shows the performance curve by the number of iterations on three representative datasets for the three task forms, GQA, NLVRv2, and RefCOCOg. We find that most performance gains occur in the first one or two iterations, after which performance plateaus and may slightly decline. Qualitative analysis shows that more iterations are beneficial for complex problems, where the initial debugging attempt often fails, so VDebugger need to iteratively refines the program in a trial-and-error manner. An example is shown in Figure 10 in the Appendix.
@@ Appendix C Experimental Details
Appendix D Qualitative Examples

### v3

@@ 1 Introduction
Complex visual reasoning is a crucial yet challenging problem that often requires compositionally synthesizing multiple reasoning steps before drawing the final conclusion. For example, to answer the visual question in Figure 1: “Do the skiers wear jackets of the same color?”, one must identify all skiers, determine the colors of their jackets, and assess whether the colors are the same. End-to-end vision-language models (VLMs) excel at individual tasks such as object detection (Li et al., 2022b) and visual instruction following (Liu et al., 2023). However, they struggle to generalize to complex tasks requiring compositional reasoning and inherently lack interpretability (Surís et al., 2023; Yüksekgönül et al., 2023; Kamath et al., 2023, 2024).
@@ 3 VDebugger Framework
where represents how likely the program is correct. is considered correct when exceeds a threshold (0.5 in this work). The critic classify the program into either correct or incorrect. Concretely, first generates a correctness token chosen from representing whether the program is correct or incorrect, so the probability assigned to token can serve as . If the token is generated, further generates the error location . Here, the location is a span within program P defined by its start and end positions, which can be a word, a line, multiple lines, or any continuous segment.
@@ 4 Training of VDebugger
and we apply mask-best sampling to at most tokens. The threshold is set as 0.9 in this work. The formal algorithm is in Alg. 2. As shown in Table 2, mask-best dramatically increases the rate at which an error is successfully injected by up to 10 times (from to ). We manually analyze and categorize 200 errors injected into GQA dataset. As shown in Figure 3, greedy sampling generates a large number of superficial errors referencing variables before their creation. In contrast, mask-best sampling produces a broader range of more complex and diverse errors.
@@ 5.1 Results
| Critic Acc. | Refiner SR | Task Performance | |||||
@@
| GQA | TallyQA | NLVRv2 | RefCOCO | RefCOCO+ | RefCOCOg | Mean | |
| CodeLlama-70b as code generator | |||||||
| No Debugging | 46.1 | 45.2 | 68.1 | 54.0 | 49.6 | 53.3 | 52.7 |
| VDebugger 7b | 48.8 (+2.7) | 47.9 (+2.7) | 68.9 (+0.8) | 57.4 (+3.4) | 50.9 (+1.3) | 56.7 (+3.4) | 55.1 (+2.4) |
| VDebugger 13b | 48.9 (+2.8) | 48.5 (+3.3) | 68.6 (+0.5) | 56.9 (+2.9) | 50.9 (+1.3) | 56.4 (+3.1) | 55.0 (+2.3) |
| DeepSeek-Coder-33B as code generator | |||||||
| No Debugging | 47.3 | 46.2 | 67.9 | 59.6 | 51.8 | 55.4 | 54.7 |
| VDebugger 7b | 48.9 (+1.6) | 46.9 (+0.7) | 67.8 (-0.1) | 60.5 (+0.9) | 52.4 (+0.6) | 57.1 (+1.7) | 55.6 (+1.9) |
| VDebugger 13b | 48.9 (+1.6) | 46.2 (+0.0) | 67.4 (-0.5) | 60.5 (+0.9) | 52.8 (+1.0) | 57.7 (+2.3) | 55.6 (+1.9) |
| GPT-3.5 as code generator | |||||||
| No Debugging | 45.9 | 43.6 | 63.9 | 60.6 | 54.0 | 57.5 | 54.3 |
| VDebugger 7b | 49.3 (+3.4) | 46.4 (+2.8) | 68.8 (+4.9) | 61.2 (+0.6) | 54.7 (+0.7) | 58.8 (+1.3) | 56.5 (+2.2) |
| VDebugger 13b | 49.6 (+3.7) | 46.4 (+2.8) | 68.7 (+4.8) | 61.3 (+0.7) | 54.1 (+0.6) | 58.8 (+1.3) | 56.5 (+2.2) |
Ablation study. We investigate the contribution of each component as shown in Table 4. Specifically, we aim to: (1) assess the individual contributions of critic and refiner components, and (2) evaluate the benefits of execution feedback. We report the critic’s binary accuracy in predicting overall program correctness, as well as the percentage of incorrect programs successfully fixed by refiner, denoted as refiner success rate. The critic demonstrates consistently strong performance, with high binary accuracy ranging from 67% to 80% across different datasets. Our manual evaluation of 59 examples from GQA shows the predicted error-inducing errors are correct in 74% of the cases. However, the refiner success rate is less reliable, varying dramatically from 10% to 57% across datasets. When enhanced with execution feedback, the critic achieves more performance gains while the benefits to refiner performance are minimal. When reflected in the final performance on the downstream tasks, execution feedback consistently brings benefits on all datasets. In general, VDebugger can reliably perform self-critique utilizing execution feedback, and the remaining challenges mainly lie in correcting the program after the errors are identified.
Performance by iteration. VDebugger can perform iterative debugging until the critic determines the program as correct. Figure 4 demonstrates the performance curve by the number of iterations on three representative datasets for the three task forms, GQA, NLVRv2, and RefCOCOg. We find that most performance gains occur in the first one or two iterations, after which performance plateaus and may slightly decline. Qualitative analysis shows that more iterations are beneficial for complex problems, where the initial debugging attempt often fails, so VDebugger need to iteratively refines the program in a trial-and-error manner. An example is shown in Figure 10 in the Appendix.
@@
Data quality. To verify the quality of automatically generated data, we manually examine 100 programs from each training set. We evaluate the proportions of incorrect programs, or "false positives", among the programs considered correct. Most datasets have relatively low false positive ratio: 16% for GQA, 13% for TallyQA, and 19% for RefCOCO. Due to the answer format, including free-form strings, numbers and bounding boxes, an exact match in the final answer ensures the program has a high probability to be correct. On the other hand, NLVRv2 dataset has a higher false positive ratio (40%) due to its binary label format. However, its effect can be mitigated by training on multiple datasets, as shown by the generalist VDebugger outperforming the specialist VDebugger on NLVRv2 dataset. While another concern over data quality is the program optimality, we observe that visual programs tend to have straightforward code structures, and thus have limited potential for algorithmic optimization. For example, 68% of the programs do not contain loop structures.
| Cost (s/it) | Compared with | Compared with | |
| No Debugging | last iteration | ||
| T = 0 | 3.08 | - | - |
| T = 1 | 4.44 | + 44% | +44% |
| T = 2 | 4.72 | + 53% | +6% |
| T = 3 | 4.81 | + 56% | +2% |
| Component | Cost (s/it) |
| Initial program generation | 1.42 |
| Program execution | 1.66 |
| Critic inference | 0.09 |
| Refiner inference | 0.08 |
@@
Computational complexity. To measure the additional computational overhead brought by the critic-refiner framework and the iterative process, we measure the computational cost by iteration step as well as a breakdown of the cost of each component in the framework. The detailed statistics are reported in Table 7. While VDebugger moderately increases the computational cost by 56% compared to the no debugging baseline, the major computational overhead arises from program execution, rather than the inference of critic and refiner models. Additionally, increasing the number of debugging iterations only marginally increases the inference cost, since most debugging is addressed within the first round.
@@ 7 Limitations
Acknowledgement
This research is based upon work supported by CISCO, U.S. DARPA ECOLE Program No. #HR00112390060, and OFFICE OF NAVAL RESEARCH Award #N00014-23-1-2780. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix C Experimental Details
| BLIP | InstructBLIP | |
| End-to-end VLMs | 43.1 | 48.3 |
| Visual programming | 45.4 | 48.4 |
| VDebugger | 48.1 | 51.1 |
Appendix D Visual Programming v.s. End-to-End VLMs
Visual programming and end-to-end VLMs are two different approaches to visual reasoning. Visual programming invokes multiple foundation VLMs through code, while end-to-end VLMs directly take an image as input and generate texts as output. Despite their seemingly different methodologies, visual programming is a complementary technique that can be combined with end-to-end VLMs to offer additional benefits. Firstly, visual programming can integrate with more powerful VLMs to further enhance performance as shown in Table 8. Secondly, despite the rapid development of end-to-end VLMs, they still have difficulty reasoning with compositional concepts such as counting and spatial relationship. Visual programming offer benefits in tasks like such as compositional reasoning, counting, and enhancing interpretability.
Appendix E Qualitative Examples